The physician ordered a creatinine clearance for an inpatient that possibly has early onset of renal problems. The nurse was instructed to collect 24 hours of urine sample with a total volume of 950 ml. The plasma creatinine was analyzed from the specimen during the early morning draw along with Complete Metabolic Panel. The result was 3.5 mg/dL. The urine chemistry result was 185 mg/dl. Calculate the creatinine clearance in ml/min.
Answer #1

Creatinine clearance

     Formula for creatinine clearance = Urine creatinine * Days urine volume /serum creatinine /1440.

                   Urine creatinine = 185 mg/dl.

                           Days urine volume = 950 ml.

                            Serum creatinine = 3.5 mg/dl.

                       As per formula     = 950 * 185 / 3.5 / 1440.

                                                            = 175750 / 3.5 / 1440.

                                                             = 50214.28 / 1440.

                                                             = 34.87 mg/dl